Please help. Can someone please show me the steps so I can understand how to do this?
Solve for a
a^2+a-42=0
Factor
72-17s+s^2

Answer by Alan3354(21588)

Solve for a
a^2+a-42=0
Factor it
(a + 7)*(a - 6) = 0
a = -7, a = 6
--------------
Factor
72-17s+s^2
= s^2 - 17s + 72 It's a good idea to put it in descending order.
----------------
The coefficents of the x terms are both 1.
So it will be:
(x ± a)*(x ± b)
The two numbers multiply to 72, a*b = 72
The 72 is positive, so they have the same sign.
The s term is negative, so a and b are negative.
--> (x - a)*(x - b)
ab = 72
a+b = -17
---------
Look at the factors of 72 and the sum:
1 and 72 = 73
2 and 36 = 38
3 and 24 = 27
4 and 18 = 22
6 and 12 = 18
8 and 9 = 17 That's it.
72-17s+s^2 = (s - 8)*(s - 9)